# Geometric quantum discord of an arbitrary two-qudit state: the exact value and general upper bounds

### v1

@@ 1 Introduction
As proved by J. Bell theoretically [1] and later experimentally by A. Aspect et al. [2], some experiments exhibit correlations between outcomes observed by different parties under joint measurements that do not agree with the classical probability model, but are describable within the quantum formalism. However, not each multipartite quantum system in all experimental scenarios gives rise to nonclassical correlations – it is a particular property of the specific quantum systems, experimental arrangements, performed quantum measurements, etc. Nonclassicality of quantum correlations is the main resource in a variety of quantum information processing tasks.
Bell nonlocality and entanglement of a multipartite quantum state are the most studied of these quantum resources and are used [3] in many quantum information protocols and algorithms. For these concepts and the quantitative and qualitative relations between them, see [4, 5, 6, 7] and references therein.
@@
Nevertheless, there are quantum states that exhibit nonclassical correlations even with small amounts of entanglement, and this led to the concept of the quantum discord [8] of a quantum state that quantifies a possible change of a quantum state under a measurement with the optimization over all possible measurements which could be performed on that state. This makes this concept conceptually rich but at the same time challenging, since it is very hard to compute it even for two qubit systems [9].
Due to the complexity of the computation of quantum discord, there are related concepts that have been proposed to quantify those quantum resources in an approximate way, although losing some of the advantages of the original concepts. This led to the appearance of some other measures quantifying nonclassicality of quantum correlations and the measurement-induced nonlocality [10] and the geometric quantum discord [11] are among them.
The geometric quantum discord is a geometric measure of quantum correlations, exhibited by a bipartite quantum state, which is defined via the distance from this state to the set $\Omega$ of all states that have the vanishing quantum discord and are referred to as classical-quantum [11] or quantum-classical [^1] In this paper, we refer to the right geometric discord instead of left discord as in [11], then the cited results will always be presented in the case of right geometric quantum discord.. This measure of quantum correlations is used in quantum phase estimation, quantum teleportation and remote state preparation protocols, see [12] and references therein.
@@
In the present article, for the definition of the geometric quantum discord we consider the Hilbert-Schmidt distance between states. In this case, the geometric quantum discord $\mathcal{D}_{g}(\rho)$ of a two-qudit state $\rho$ on a complex Hilbert space $\mathcal{H}_{d_{1}}\otimes\mathcal{H}_{d_{2}}$ is defined by the relation
$\mathcal{D}_{g}(\rho):=\min_{\chi\in\Omega}||\rho-\chi||_{2}^{2},$ (1)
where $||\cdot||_{2}$ denotes the Hilbert-Schmidt norm. Possible other definitions [13, 14, 15] of the geometric quantum discord are based on a different than in (1) distance between quantum states.
@@
Though the optimization problem involved into the computation of the geometric quantum discord is much simpler than that for the quantum discord, the geometric quantum discord of a bipartite state has been computed explicitly only in some particular cases, namely, for a general two-qubit state [11], a general qubit-qudit state [16] and some special families of two-qudit states [16, 17].
@@
However, to our knowledge, the explicit exact analytical value of the geometric quantum discord for a general two-qudit state of an arbitrary dimension has not been reported in the literature – only its lower bounds, see in [16, 17, 18, 19, 20].
@@
In the present paper, for an arbitrary two-qudit state $\rho$ on $\mathcal{H}_{d_{1}}\otimes\mathcal{H}_{d_{2}},$ pure or mixed, we find the explicit exact analytical value of its geometric quantum discord (1) via parameters of its correlation matrix and the Bloch vectors of its reduced states. This new result: (i) reduces in a two-qubit case to the well-known two-qubit expression of Dakić et al. [11]; (ii) proves that the lower bound found by Rana and Parashar in [16] is attained on each two-qudit state; (iii) allows us to find the new general upper bounds on the geometric quantum discord of an arbitrary two-qudit state in terms of its Hilbert space characteristics and in case of a pure two-qudit state – via its concurrence.
@@ 2 Preliminaries: general Bloch vectors formalism
$T_{\rho_{d_{1}\times d_{2}}}^{(ij)}=\mathrm{tr}\left[\rho_{d_{1}\times d_{2}}\left(\Upsilon_{d_{1}}^{(i)}\otimes\Upsilon_{d_{2}}^{(j)}\right)\right],\text{\ \ }i=1,...,d_{1}^{2}-1,\text{ \ \ }j=1,...,d_{2}^{2}-1,$ (15)
@@
$\left\|T_{\rho_{d\times d}}\right\|_{0}$ $:=\sup_{\left\|n\right\|=1}\left\|T_{\rho_{d\times d}}n\right\|_{\mathbb{R}^{d^{2}-1}}=\sqrt{d-1}\sup_{\left\|n\right\|=1}\left\|T_{\rho_{d\times d}}(\frac{n}{\sqrt{d-1}})\right\|_{\mathbb{R}^{d^{2}-1}}$ (20)
@@ 3 Geometric quantum discord
Let [^2] See footnote 1.
@@
$+\frac{1}{4}\sum_{\begin{subarray}{c}i=1,...,d_{1},\\ j=1,...,d_{2}\end{subarray}}\left(T_{\rho}^{(ij)}-2\sqrt{\frac{(d_{1}-1)(d_{2}-1)}{d_{1}d_{2}}}\sum_{k}\alpha_{k}x_{k}^{(i)}y_{k}^{(j)}\right)\Upsilon_{d_{1}}^{(i)}\otimes\Upsilon_{d_{2}}^{(j)},$ (33)
@@
$\sum_{k=1}^{d_{2}}y_{k}=0,\text{ }\left\|y_{k}\right\|_{\mathbb{R}^{d_{2}^{2}-1}}=1,\text{ \ \ }y_{k}\cdot y_{k_{1}}=-\frac{1}{d_{2}-1}$ (35)
@@ 4 Upper and lower bounds
$\mathcal{D}_{g}(\rho)$ $\leq\frac{d_{2}-1}{d_{2}}\left[1-\frac{\left\|r_{2}\right\|^{2}}{d_{2}+1}\right]$ (57)
@@
$=\frac{d_{1}d_{2}-1}{d_{1}(d_{2}+1)}-\frac{1}{d_{2}+1}\cdot\frac{d_{1}-1}{d_{1}}\left\|r_{1}\right\|^{2}.$
@@
$\sum_{n=d_{2}}^{d_{2}^{2}-1}\eta_{n}$ $\leq\frac{d_{1}d_{2}-1}{d_{1}(d_{2}+1)}-\frac{1}{d_{2}+1}\left(\frac{1}{d_{2}}-\frac{1}{d_{1}}+\frac{d_{2}-1}{d_{2}}\left\|r_{2}\right\|^{2}\right)$ (62)
$=\frac{d_{2}-1}{d_{2}}\left[1-\frac{\left\|r_{2}\right\|^{2}}{d_{2}+1}\right].$
@@
where $r_{j}^{(k)}$ are the Bloch vectors of states $\rho_{j}^{(k)},$ $j=1,2,$ given by (5) with norms $||r_{1}^{(k)}||,$ $||r_{2}^{(k)}||$ $\leq 1.$ If a separable state is pure, then $||r_{1}||=$ $||r_{2}||$ $=1$ and the positive Hermitian operator (48) is given by
$G(\rho_{sep}^{(pure)})$ $=\frac{d-1}{d^{2}}|r_{2}\rangle\langle r_{2}|\text{ }+\text{ }\left(\frac{d-1}{d}\right)^{2}\ \left\|r_{1}\right\|^{2}|r_{2}\rangle\langle r_{2}|$ (75)
@@ 6 Appendix A
the rank of every projection $\Pi_{\Omega_{y}}$ in Proposition 2 is equal to $\left(d-1\right).$ Note that an orthogonal projection $\Pi_{\Omega_{y}}$ has eigenvalue $1$ of multiplicity $(d-1)$ and eigenvalue 00 of multiplicity $d(d-1).$ Let
Denote by $e_{n},$ $n=1,...,N,$ mutually orthogonal unit eigenvectors of a positive Hermitian operator $A,$ corresponding to its eigenvalues $\zeta_{1}\geq\zeta_{2}\geq\cdots\geq\zeta_{N}\geq 0,$ listed in the decreasing order with the corresponding algebraic multiplicities,.and by $g_{m}\in\mathcal{H}_{N},$ $k=1,...,N,$ mutually orthogonal eigenvectors of projection $\Pi_{\Omega_{y}},$ where the first $(d-1)$ eigenvectors correspond to eigenvalue $1$. By the spectral theorem,
@@ 6.1 Examples
$=\sum_{n=1}^{d-1}|e_{n}\rangle\langle e_{n}|\text{ }+\text{ }\frac{1}{d}\sum_{n,n^{\prime}=1}^{d-1}\left(\sum_{k=1}^{d}\alpha_{kn}\alpha_{kn^{\prime}}\right)|e_{n}\rangle\langle e_{n^{\prime}}|$
@@
$\sum_{k=2}^{d}\alpha_{kn}$ $=0,\text{ \ \ }n=2,...,(d-1),\text{ \ \ }\sum_{d=2}^{d-1}\alpha_{k_{1}n}\alpha_{k_{2}n}=-1,\text{ \ \ }k_{1}\neq k_{2}\in\{2,...,d\},$
@@
In each vector $y_{k\geq 3}^{\prime},$ the number $Z_{d}(+)$ of positive $\alpha_{kn}=+1$ is equal to $Z_{d}(+)=\frac{d-3}{2}$ and the number $Z_{d}(-)$ of $\alpha_{kn}=-1,k\neq 1,$ is equal to $\frac{d-1}{2}.$ For these vectors,
$\sum_{k=1}^{d}y_{k}^{\prime}=0,\text{ \ \ \ }y_{k_{1}}^{\prime}\cdot y_{k_{2}}^{\prime}=-\frac{1}{d-1},$ (A13)
@@
$\sum_{k}|y_{k}^{\prime}\rangle\langle y_{k}^{\prime}|\text{ }$ $=|e_{1}\rangle\langle e_{1}|\text{ }+\text{ }\frac{d-1}{(d-1)^{2}}|e_{1}\rangle\langle e_{1}|$ (A14)
@@
    $d=5$: then in each $y_{k\geq 3}^{\prime}$, the number $Z_{5}(+)$ of $\alpha_{kn}=1$ is equal to $1$ and the number $Z_{5}(-)_{-}$ of $\alpha_{kn}=-1$ is equal to $2$ The vectors

### v2

@@ 1 Introduction
As shown by J. Bell theoretically [1] and later experimentally by A. Aspect et al. [2], the probabilistic description of a quantum correlation scenario does not, in general, agree with the classical probability model. Nonclassicality of quantum correlations is one of the main resources for many quantum information processing tasks. Among these quantum resources, Bell nonlocality and entanglement are the most studied, see [3, 4, 5, 6] and references therein for the quantitative and qualitative relations between them.
@@
Nevertheless, there are quantum states that exhibit nonclassical correlations even without entanglement, and this led to the notion of the quantum discord [7], which is conceptually rich, however, it is very hard to calculate it even for a two-qubit state [8].
Due to the complexity [9] of computation of the quantum discord, there were also introduced related concepts, like the measurement-induced nonlocality [10] and the geometric quantum discord [11].
The geometric quantum discord is a geometric measure of quantum correlations of a bipartite quantum state, which is defined via the distance from this state to the set $\Omega$ of all states with the vanishing quantum discord [11].
@@
In the present article, the geometric quantum discord $\mathcal{D}_{g}(\rho)$ of a two-qudit state $\rho$ on $\mathcal{H}_{d_{1}}\otimes\mathcal{H}_{d_{2}}$ is defined via the Hilbert-Schmidt distance $||\cdot||_{2}$ between states:
$\mathcal{D}_{g}(\rho)=\min_{\chi\in\Omega}||\rho-\chi||_{2}^{2}=\min_{\chi\in\Omega}\mathrm{tr}[(\rho-\chi)^{2}]\ .$ (1)
In other definitions [12, 13, 14] of the geometric quantum discord, different than in (1) distances are used.
@@
Though the optimization problem for the computation of the geometric quantum discord of a bipartite state is much simpler than that for the quantum discord, its exact value has been explicitly computed only in some particular cases, namely, for a general two-qubit state [11], a general qubit-qudit state [15] and some special families of two-qudit states [15, 16].
@@
However, to our knowledge, for a general two-qudit state of an arbitrary dimension, the explicit exact analytical value of the geometric quantum discord has not been reported in the literature – only its lower bounds [15, 16, 17, 18, 19].
@@
In the present paper, for an arbitrary two-qudit state $\rho$ on $\mathcal{H}_{d_{1}}\otimes\mathcal{H}_{d_{2}},$ pure or mixed, we find in the explicit form the exact analytical value of its geometric quantum discord (1). This new rigorously proved general result indicates that the lower bound found in [15] is attained on each two-qudit state and includes only as particular cases the exact results for a general two-qubit state [11], a general qubit-qudit state [15] and some special families of two-qudit states [15, 16]. It also allows us to find the new general upper bounds on the geometric quantum discord of an arbitrary two-qudit state in terms of its Hilbert space characteristics and in case of a pure two-qudit state – via its concurrence.
@@ 2 Preliminaries: general Bloch vectors formalism
$T_{\rho_{d_{1}\times d_{2}}}^{(ij)}=\mathrm{tr}\left[\rho_{d_{1}\times d_{2}}\left(\Upsilon_{d_{1}}^{(i)}\otimes\Upsilon_{d_{2}}^{(j)}\right)\right],\text{\ \ }i=1,...,(d_{1}^{2}-1),\text{ \ \ }j=1,...,(d_{2}^{2}-1),$ (15)
@@
$\left\|T_{\rho_{d\times d}}\right\|_{0}$ $:=\sup_{\left\|n\right\|=1}\left\|T_{\rho_{d\times d}}n\right\|_{\mathbb{R}^{d^{2}-1}}=\sqrt{d-1}\sup_{\left\|n\right\|=1}\left\|T_{\rho_{d\times d}}\left(\frac{n}{\sqrt{d-1}}\right)\right\|_{\mathbb{R}^{d^{2}-1}}$ (20)
@@ 3 Geometric quantum discord
Let [^1] In this paper, we refer to the right geometric discord instead of left discord as in [11].
@@
$+\frac{1}{4}\sum_{\begin{subarray}{c}i=1,...,d_{1},\\ j=1,...,d_{2}\end{subarray}}\left(T_{\rho}^{(ij)}-2\sqrt{\frac{(d_{1}-1)(d_{2}-1)}{d_{1}d_{2}}}\sum_{k}\alpha_{k}x_{k}^{(i)}y_{k}^{(j)}\right)\Upsilon_{d_{1}}^{(i)}\otimes\Upsilon_{d_{2}}^{(j)},$
@@
$\sum_{k=1}^{d_{2}}y_{k}=0,\text{ }\left\|y_{k}\right\|_{\mathbb{R}^{d_{2}^{2}-1}}=1,\text{ \ \ }y_{k_{1}}\cdot y_{k_{2}}=-\frac{1}{d_{2}-1},\text{ \ \ \ }\forall k_{1}\neq k_{2}.$ (32)
@@ 4 Upper and lower bounds
$\mathcal{D}_{g}(\rho)$ $\leq\frac{d_{2}-1}{d_{2}}\left[1-\frac{\left\|r_{2}\right\|^{2}_{\mathbb{R}^{d_{2}^{2}-1}}}{d_{2}+1}\right]$ (54)
@@
$=\frac{d_{1}d_{2}-1}{d_{1}(d_{2}+1)}-\frac{1}{d_{2}+1}\cdot\frac{d_{1}-1}{d_{1}}\left\|r_{1}\right\|_{\mathbb{R}^{d_{1}^{2}-1}}^{2}.$
@@
$\sum_{n=d_{2}}^{d_{2}^{2}-1}\eta_{n}$ $\leq\frac{d_{1}d_{2}-1}{d_{1}(d_{2}+1)}-\frac{1}{d_{2}+1}\left(\frac{1}{d_{2}}-\frac{1}{d_{1}}+\frac{d_{2}-1}{d_{2}}\left\|r_{2}\right\|^{2}_{\mathbb{R}^{d_{2}^{2}-1}}\right)$ (59)
$=\frac{d_{2}-1}{d_{2}}\left[1-\frac{\left\|r_{2}\right\|^{2}_{\mathbb{R}^{d_{2}^{2}-1}}}{d_{2}+1}\right].$
@@
where $r_{j}^{(k)}$ are the Bloch vectors of states $\rho_{j}^{(k)},$ $j=1,2,$ given by (5) with norms $||r_{1}^{(k)}||_{\mathbb{R}^{d^{2}-1}},$ $||r_{2}^{(k)}||_{\mathbb{R}^{d^{2}-1}}$ $\leq 1.$ If a separable state is pure, then $||r_{1}||_{\mathbb{R}^{d^{2}-1}}=$ $||r_{2}||_{\mathbb{R}^{d^{2}-1}}$ $=1$ and the positive Hermitian operator (45) is given by
$G(\rho_{sep}^{(pure)})$ $=\frac{d-1}{d^{2}}|r_{2}\rangle\langle r_{2}|\text{ }+\text{ }\left(\frac{d-1}{d}\right)^{2}\ \left\|r_{1}\right\|_{\mathbb{R}^{d^{2}-1}}^{2}|r_{2}\rangle\langle r_{2}|$ (72)
@@ 6 Appendix A
the rank of every projection $\Pi_{\Omega_{y}}$ in Proposition 2 is equal to $\left(d-1\right).$ Note that an orthogonal projection $\Pi_{\Omega_{y}}$ has eigenvalue $1$ of multiplicity $(d-1)$ and eigenvalue 00 of multiplicity $d(d-1).$
Denote by $e_{n},$ $n=1,...,N,$ where $N=d^{2}-1,$ mutually orthogonal unit eigenvectors of a positive Hermitian operator $A,$ corresponding to its eigenvalues $\zeta_{1}\geq\zeta_{2}\geq\cdots\geq\zeta_{N}\geq 0,$ listed in the decreasing order with the corresponding algebraic multiplicities, and by $g_{n},n=1,...,N,$ mutually orthogonal eigenvectors of projection $\Pi_{\Omega_{y}},$ where the first $(d-1)$ eigenvectors correspond to eigenvalue $1$. By the spectral theorem,
@@ 6.1 Examples
$=\sum_{n=1}^{d-1}|e_{n}\rangle\langle e_{n}|\text{ }+\text{ }\frac{1}{d}\sum_{n\neq n^{\prime}\in\{1,...,d-1\}}\left(\sum_{k=1}^{d}\alpha_{kn}\alpha_{kn^{\prime}}\right)|e_{n}\rangle\langle e_{n^{\prime}}|$
@@
$\sum_{k=2}^{d}\alpha_{kn}$ $=0,\text{ \ \ }n=2,...,(d-1),\text{ \ \ }\sum_{n=2}^{d-1}\alpha_{k_{1}n}\alpha_{k_{2}n}=-1,\text{ \ \ }k_{1}\neq k_{2}\in\{2,...,d\},$
@@
In each vector $y_{k\geq 3}^{\prime},$ the number $Z_{d}(+)$ of positive $\alpha_{kn}=+1$ is equal to $Z_{d}(+)=\frac{d-3}{2}$ and the number $Z_{d}(-)$ of $\alpha_{kn}=-1,k\neq 1,$ is equal to $\frac{d-1}{2}.$ For these unit vectors,
$\sum_{k=1}^{d}y_{k}^{\prime}=0,\text{ \ \ \ }y_{k_{1}}^{\prime}\cdot y_{k_{2}}^{\prime}=-\frac{1}{d-1},\ \ k_{1}\neq k_{2}\ .$ (A13)
@@
$\sum_{k=1}^{d}|y_{k}^{\prime}\rangle\langle y_{k}^{\prime}|\text{ }$ $=|e_{1}\rangle\langle e_{1}|\text{ }+\text{ }\frac{d-1}{(d-1)^{2}}|e_{1}\rangle\langle e_{1}|$ (A14)
@@
    $d=5$: then in each $y_{k\geq 3}^{\prime}$, the number $Z_{5}(+)$ of $\alpha_{kn}=1$ is equal to $1$ and the number $Z_{5}(-)$ of $\alpha_{kn}=-1$ is equal to $2$ The vectors